OpenStudy (anonymous):

I can't figure out how to solve these four, I know the answers but don't know the steps. Thanks 2x-3/4 +5=3x (t+5)/8 - (t-2)/2= 1/3 (x-5)/4 + (3-2x)/3 > -1 2y-3/2 + 3y-1/5 < y-1

OpenStudy (tkhunny):

Please show your best efforts on the first one.

OpenStudy (anonymous):

Do you mean: \(\dfrac{2x-3}{4} + 5 = 3x\) or \(2x-\dfrac{3}{4}+5=3x\)?

OpenStudy (anonymous):

The first one u typed , but I solve the first one I figured it out, I just need help one the last three. Do I tried cross multiplying but I don't get the right answer.

OpenStudy (texaschic101):

I think you multiply by the common denominator

OpenStudy (anonymous):

\(\dfrac{t+5}{8} - \dfrac{t-2}{2}= \dfrac{1}{3}\) You can only cross multiply if there is one term on each side. In this case, you can combine the terms on the left side to make one fraction and then cross multiply.

OpenStudy (anonymous):

Or you can multiply through the entire equation by a common denominator.

OpenStudy (anonymous):

So the common denominator would be 24 for all three?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So I should get 3t-5/24 - 12t-24/24 = 8/24

OpenStudy (anonymous):

The point of multiplying by the common denominator is to eliminate the fractions. I'll demonstrate on the first term.

OpenStudy (anonymous):

\(\dfrac{\cancel{24}(3)}{1} \times \dfrac{t-5}{\cancel{8}} = 3t+5\)

OpenStudy (anonymous):

oh ok so it will be 3t+5-12t-24=8 and solve

OpenStudy (anonymous):

Remember to distribute the negative in the second term.

OpenStudy (anonymous):

Thank you so much! I got 31/9. And for the other two I just do the same?

OpenStudy (anonymous):

Close. Did you actually get 31/9 or did you mistype?

OpenStudy (anonymous):

Yeah I mistyped, but I got that. thank you!

OpenStudy (anonymous):

Good job! And yes, the other questions will be done the same way.

OpenStudy (texaschic101):

I have a question.....after you multiply by 24, do you get 3(t + 5) or just 3t + 5

OpenStudy (anonymous):

@$%&^! Your right, I forgot to distribute.

OpenStudy (texaschic101):

I wasn't sure......thats why I asked

OpenStudy (texaschic101):

thanks for explaining it

OpenStudy (anonymous):

@rrrgirls1 did get it right though.